4. (Heath 4.3) Given an approximate eigenvector x of A, what is the best estimate (in the least squares sense) for the corresponding eigenvalue?
Answer #1

Rayleigh Quotient

Given an approximate eigenvector x of A, the Rayleigh Quotient is the best estimate.

Consider the eigenvalue equation xλ = Ax where x is our current best eigenvector guess and λ is unknown. the coefficient matrix is the n × 1 matrix x. The normal equations say that the least squares answer is the solution of x^ T xλ = x ^T Ax, or λ = (x^T Ax) / (x ^T x) . λ is known as the Rayleigh Quotient.

------------------------

thank you